OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y =1/3x-3 A) y=3x-14 B) y= 1/3x+2 C) y=-3x+22 D) y=-1/3x+8

OpenStudy (anonymous):

since both lines are parallel what do u know?

OpenStudy (anonymous):

well im not really sure how to do these..

OpenStudy (anonymous):

okay since the line is parallel to y =1/3x-3, then 1/3 is the gradient since y=mx+c where m is gradient, therefore the answer straight away is B where gradient of B is also 1/3

OpenStudy (anonymous):

Ohhhh! Thank you so much :)

OpenStudy (anonymous):

@LauKang could you help me with a couple others???? :)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (2, -3) and is perpendicular to y = 2/3x+3 A) y=-3/2x B) y=3/2x-6 C) y=-2/3-5/3 D) y=2/3x-13/3

OpenStudy (anonymous):

okay for perpendicular lines note this

OpenStudy (anonymous):

if line A is perpendicular to Line B, gradient of A X gradient of B = -1

OpenStudy (anonymous):